# Real time observation of glass-like carbon formation from SU-8 using X-ray and ultraviolet photoelectron spectroscopy

### v1

@@ 3 Results & Discussion
Figures 4 and 5 shows detailed XPS spectra at binding energies in the C 1s (280 - 292 eV) and O1s (528 - 542 eV) regions respectively. Selected C 1s and O 1s spectra are shown which have been fitted using Voigt functions (a convolution of a Gaussian and Lorentzian line-shape). A minimum number of peaks were selected for each spectrum, using physically meaningful variables for Gaussian and Lorentzian contributions related to the sample and instrumentation used. The fitting process was iterated to minimise residuals. An asymmetry parameter was introduced to reflect the changing conductivity from semiconducting to conducting and therefore, flexibility in the asymmetry fitting was allowed for the C 1s feature. The relative Gaussian and Lorentzian ratios were kept consistent throughout this whilst allowing for broadening due to measurements at high temperatures.
@@
The C 1s core level emission spectra (Figure 4) were fitted with a combination of peaks that are associated broadly with C-C, C-H and C-O bonding within the SU-8 molecule and, at higher binding energy, contaminant carbon atoms not associated with the SU-8 sample. Due to the instrumental energy resolution and the semiconducting nature of the sample, it is not possible to ascribe precise bonding environments to the core level components [26, 27]. The highest binding energy component, present at 100 C is identified due to contaminant carbon species that are not associated with the SU-8. These are desorbed as temperature increases, with no contributions at temperatures above 300 C. The core level line-shape and position is not reproducible for different samples as-loaded, following ex-situ preparation. However, the XPS spectra are consistent for all samples following annealing to 300 C in UHV which suggests clean, contamination-free SU-8. Fitting of the lower-temperature surfaces is thus more subject to variation and less representative of the SU-8 molecule. The C 1s spectrum at 400 C can be adequately fitted with two components as shown in Figure 4. The lower intensity component (C-O group 1) is around 10% of the more intense component (C-C group 1) and this is similar to the overall oxygen concentration at this temperature (Figure 3). This suggests that the lower intensity component is due to carbon-oxygen bonds with other carbon species contributing to the more intense peak. The component at a binding energy of 286.7 eV (C-O group 1 in Figure 4) represents a number of possible C-O bonding environments in the matrix. C-O Group 1 is present at 300 C and it is not present at the higher temperatures. This is interpreted as the desorption of oxygen from the SU-8 as the glass phase is formed. The main C 1s component at a binding energy of 285 eV is due carbon atoms bound to other carbon atoms and to hydrogen atoms within the SU-8 ( C-C Group 1 in Figure 4). Within the energy resolution of the instrument, it is not possible to unambiguously identify the different C-C bonding sites nor C-H bonding where the carbon atoms have similar symmetry. The binding energy of these components are close together for other carbon containing materials [28, 29] and, in a combined experimental and theoretical study of ether and epoxy containing material, the binding energy for these two sites are supposed to be the same [30]. For semiconductors, core level binding energy can also be affected by local band bending and changes in band gap that can further contribute to the overall peak broadening [31]. At lower temperatures, a further component at higher binding energy (C-C group 2) is required to obtain a reasonable fit. The relative intensity of this component is preparation-
@@
Since the identification of precise C-O bonding is not possible from the C 1s emission spectra even for O-terminated, single crystal diamond, [38] the O 1s core level spectra have been fitted as shown in Figure 5 for temperatures up to 650 C. Each spectrum has been fitted with three peaks (referred to as C-O groups in Figure 3) with an additional peak attributed to O-O bonds at 540 eV.
@@
As shown in Figure 1A, there are two oxygen environments in the SU-8 molecule : one attached to an aromatic ring in a R-O-R ether configuration and the other in an epoxy configuration [18, 20]. Each of these could have a distinct binding energy and thermal degradation mechanisms [25]. At 300 C, there are two components in the fitted curve around 533 eV that are identified as the ether and epoxy sites in the SU-8 molecule. The binding energy difference is small due to the similarity in the oxygen environment where each oxygen atom is singly-bonded to one carbon atom. The intensity of the epoxy oxygen component is lower, suggesting that a preferential desorption has commenced at 300 C. By 550 C, only the ether oxygen component remains, and this disappears by 650 C as these oxygen atoms are desorbed from the molecule. At 400 C, a new component in the O 1s core level emission spectra at lower binding energy is required to obtain a reasonable fit (C-O group 3). Components at these binding energies are usually identified as due to double-bonded oxygen atoms in a ketone configuration. The binding energy of oxygen atoms in groups such as epoxy and ether can vary considerably for different molecules [30] and these authors also point out that charging can also lead to ambiguous assignment of sp3 carbon. Comparison with other methods such as IR and Raman can confirm the presence of different oxygen species [28, 29, 30]. In the present study, the fitting parameter for the O 1s and C 1s core levels are consistent with the literature and the accepted model for Su-8 molecule discussed in Figure 1. The peak located at 540.4 eV is attributed to gaseous oxygen (O2) [39] and is due to oxygen desorbing from the structure at a slow rate due to the reduced porosity of glass-like carbon. The C-O group 3 therefore probably arises from the conversion of epoxy oxygen in the outer arm to ketone like oxygen by the removal of a CH2 group. At 550 C, the ketone oxygen is desorbed and the aromatic rings start to fuse into the graphitic structure shown in Figure 1. At temperatures below 300 C there is likely to be a contribution to the O 1s core level spectra from oxygen contaminants or fragments and not due to the SU-8 molecule. The 100 C spectra in Figure 3 has been fitted using two components that are likely to contain contaminant and SU-8 oxygen sites.
@@
The binding energy of each fitted peak for both the C 1s and O 1s regions is shown in Figure 6. In the C 1s spectra, contamination carbon is only present at low temperatures and is removed at temperatures above 100 C. The position of the peak associated with C-O bonds in the C 1s spectra is largely unchanged, and this can be also seen in the O 1s spectra with all three C-O groups (Figure 6B), indicating little change in binding energy throughout pyrolysis. The main peak, dominated by C-C bonding, in the C 1s spectra exhibits the most considerable change throughout the pyrolysis process. An initial decrease in binding energy between 100 - 200 C is followed, at higher temperatures, by a decrease in binding energy from 285.0 eV to 284.2 eV. Shifts in core levels can be due to a variety of physical and chemical changes [31] whilst changes in the relative component intensities can be used to inform the processes and mechanisms for these shifts. The final peak position for the main C 1s component (C-C Group 1) at a binding energy of 284.2 eV is indicative of a carbon structure similar to that of sp2 carbon [26, 40, 41]. This suggests that whilst the elimination of oxygen during carbonisation has little effect on the structure, the elimination of hydrogen at temperature above 500 C is key to development of an sp2 network indicative of glass-like carbon.
@@
The shift in binding energy thus indicates the transition from a semiconducting to conducting material between 500 C and 700 C. The intensity of the O 1s peak is low at these temperatures and hence it is not possible to determine with accuracy the binding energy of oxygen components in this temperature range. For all measurements of SU-8 at 300 C and its conversion into a graphitic network at higher temperatures,are not affected by charging due to the photoemission process. Conductivity measurements [42] on this material yield a centre to edge resistance of 325k Ohms at 700 C and 4k Ohms at 750 C. From top to bottom the corresponding resistances are 0.46 ohms to 0.03 ohms between 700 C and 750 C respectively. For typical values of photocurrent generated in the XPS spectrometer, significant charging shift would only be observed for a resistance of greater than 100M Ohms. The initial large shift in binding energy in the C 1s and O 1s suggests that the presence of contaminants affects the bulk electronic nature of the measured sample surface, confirming that a true measurement of SU-8 is not achieved until 300 C.
@@
The C-C peak also exhibits an increase in asymmetry as a result of being heated to higher pyrolysis temperature. Figure 7 shows the variation of asymmetry factor (, as defined by the Voigt fit) with pyrolysis temperature. A sharp increase in asymmetry occurs at 550 C that continues to increase to a value of 0.95 at a pyrolysis temperature of 1000 C.
@@
Asymmetry is an important consideration in XPS analysis of conducting materials. Asymmetric peaks are due to a variety of factors such as vibrational modes, multi electron excitations and electron - hole pair generation. In carbon structures such as graphite, the asymmetry can be ascribed to electrons excited into the conduction band during the photoemission process, causing loss of energy for the core level photoelectrons. As the sample becomes more conductive, the change in asymmetry can be correlated with changing conductivity [43, 44, 45].
@@
Figure 10A shows in more detail the valence band maximum of SU-8 as a function of temperature with spectra at selected temperatures shown in Figure 10B with a focus on photoelectrons emitted from near the Fermi Energy at 21.2 eV. The spectra were recorded with a small negative sample bias to enable measurement of the SECO, this value has been subtracted from the UPS spectra shown in Figures 9 and 10. As the temperature increases, a broad peak appears at 17 eV, that is most prominent at 400 C and disappears in the temperature range of 700 - 750 C. This coincides with the disappearance of C-O group 2 beyond 300 C (Figure 3) and the saturation of C-C peak position and is thus likely to be due to surface oxygen species (Figure 4). As the temperature approaches 1000 C, the VBM shows increasing emission at the Fermi energy (zero binding energy) at 21.2 eV (inset Figure 10B) that indicates the full transformation of insulating SU-8 to fully conducting glass-like carbon reported in the literature [15].

### v2

@@ 3 Results & Discussion
Figures 4 and 5 shows detailed XPS spectra at binding energies in the C 1s (280 - 292 eV) and O1s (528 - 542 eV) regions respectively. Selected C 1s and O 1s spectra are shown which have been fitted using Voigt functions (a convolution of a Gaussian and Lorentzian line-shape). A minimum number of peaks were selected for each spectrum, using physically meaningful variables for Gaussian and Lorentzian contributions related to the sample and instrumentation used. The fitting process was iterated to minimise residuals. An asymmetry parameter was introduced to reflect the changing conductivity from semiconducting to conducting and therefore, flexibility in the asymmetry fitting was allowed for the C 1s feature. The relative Gaussian and Lorentzian ratios were kept consistent throughout whilst allowing for broadening due to measurements at high temperatures.
@@
The C 1s core level emission spectra (Figure 4) were fitted with a combination of peaks that are associated broadly with C-C, C-H and C-O bonding within the SU-8 molecule and, at higher binding energy, contaminant carbon atoms not associated with the SU-8 sample. Due to the instrumental energy resolution and the semiconducting nature of the sample, it is not possible to ascribe precise bonding environments to the core level components [26, 27]. The highest binding energy component (C3), present at 100 C is identified due to contaminant carbon species that are not associated with the SU-8 molecule. These are desorbed as temperature increases, with no contributions at temperatures above 300 C. The core level line-shape and position is not reproducible for different samples as-loaded, following ex-situ preparation. However, the XPS spectra are consistent for all samples following annealing to 300 C in UHV which suggests clean, contamination-free SU-8. Fitting of the lower-temperature surfaces is thus more subject to variation and less representative of the SU-8 molecule. The C 1s spectrum at 400 C can be adequately fitted with two components as shown in Figure 4. The lower intensity component (C2) is around 10% of the more intense component (C1) and this is similar to the overall oxygen concentration at this temperature (Figure 3). The component at a binding energy of 286.7 eV (C2 in Figure 4) represents a number of possible C-O bonding environments in the matrix. C2 is present at 300 C and it is not present at the higher temperatures. This is interpreted as the desorption of oxygen from the SU-8 as the glass phase is formed. The main C 1s component at a binding energy of 285 eV is due to carbon atoms bound to other carbon atoms and to hydrogen atoms within the SU-8 molecule ( C1 in Figure 4). Within the energy resolution of the instrument, it is not possible to unambiguously identify the different C-C bonding sites nor C-H bonding where the carbon atoms have similar symmetry. The binding energy of these components are close together for other carbon containing materials [28, 29] and, in a combined experimental and theoretical study of ether and epoxy containing material, the binding energy for these two sites are reported to be the same [30]. For semiconductors, core level binding energy can also be affected by local band bending and changes in band gap that can further contribute to the overall peak position and peak broadening [31]. At lower temperatures, a further component at higher binding energy (C3) is required to obtain a reasonable fit. The relative intensity of this component is preparation-
@@
Since the identification of precise C-O bonding is not possible from the C 1s emission spectra even for O-terminated, single crystal diamond, [38] the O 1s core level spectra have been fitted as shown in Figure 5 for temperatures up to 650 C. Each spectrum has been fitted with three peaks (O1, O2 and O3 in Figure 3) with an additional peak, O4, not associated with the SU-8 molecule, at 540 eV.
@@
As shown in Figure 1A, there are two oxygen environments in the SU-8 molecule : one attached to an aromatic ring in a R-O-R ether configuration (O1) and the other in an epoxy configuration (O2) [18, 20]. Each of these could have a distinct binding energy and thermal degradation mechanisms [25]. At 300 C, the fitted curve only requires the two components, O1 and O2, corresponding to the ether and epoxy sites in the SU-8 molecule. The binding energy difference is small due to the similarity in the oxygen environment where each oxygen atom is singly-bonded to two carbon atom. The intensity of the epoxy oxygen component is lower, suggesting that a preferential desorption has commenced at 300 C. By 550 C, only the ether oxygen component remains, and this disappears by 650 C as these oxygen atoms are desorbed from the molecule. At 400 C, a new component in the O 1s core level emission spectra at lower binding energy is required to obtain a reasonable fit (O3). Components at these binding energies are usually identified as due to double-bonded oxygen atoms in a ketone configuration [39]. The binding energy of oxygen atoms in groups such as epoxy and ether can vary considerably for different molecules [30] and these authors also point out that charging can also lead to ambiguous assignment of sp3 carbon. Comparison with other methods such as IR and Raman can confirm the presence of different oxygen species [28, 29, 30]. In the present study, the fitting parameters for the O 1s and C 1s core levels are consistent with the literature and the accepted model for SU-8 molecule discussed in Figure 1a. The peak located at 540.4 eV is not commonly observed for solids containing carbon and oxygen, but is close in energy to gaseous, molecular oxygen (O2) observed in near-ambient XPS [31, 40, 41, 42, 43, 44]. Oxygen is released during the transition from semiconducting to conducting phases and this oxygen desorbing from the structure at a slow rate may be detectable due to the porosity of the material. The O3 component is therefore likely to arise from the conversion of epoxy oxygen in the outer arm to ketone-like oxygen by the removal of a CH2 group. At 550 C, the ketone oxygen is desorbed and the aromatic rings start to fuse into the graphitic structure shown in Figure 1b. At temperatures below 300 C there is likely to be a contribution to the O 1s core level spectra from oxygen contaminants or fragments and not due to the SU-8 molecule. The 100 C spectra in Figure 3 has been fitted using two components that are likely to contain contaminant and SU-8 oxygen sites.
@@
The temperature-dependence of the binding energy of each fitted peak for both the C 1s and O 1s regions is shown in Figure 6. In the C 1s spectra, contamination carbon is only present at low temperatures and is removed at temperatures above 100 C. The position of the peak associated with C-O bonds in the C 1s spectra is largely unchanged, and this can be also seen in the O 1s spectra with all three O groups (Figure 6B), indicating little change in binding energy throughout pyrolysis. The main peak, dominated by C-C bonding, in the C 1s spectra exhibits the most considerable change throughout the pyrolysis process. An initial decrease in binding energy between 100 - 300 C is followed, at temperatures above 500 C, by a further decrease in binding energy from 285.0 eV to 284.2 eV. Shifts in core levels can be due to a variety of physical and chemical changes [31] whilst changes in the relative component intensities can be used to inform the processes and mechanisms for these shifts. The final peak position for the main C 1s component (C1) at a binding energy of 284.2 eV is indicative of a carbon structure similar to that of sp2 carbon [26, 45, 46]. This suggests that whilst the elimination of oxygen during carbonisation has little effect on the structure, the elimination of hydrogen at temperature above 500 C is key to development of an sp2 network indicative of glass-like carbon.
@@
The shift in binding energy thus indicates the transition from a semiconducting to conducting material between 500 C and 700 C. The intensity of the O 1s peak is low at these temperatures and hence it is not possible to determine with accuracy the binding energy of oxygen components in this temperature range. For all measurements of SU-8 at 300 C and its conversion into a graphitic network at higher temperatures, are not affected by charging due to the photoemission process. Conductivity measurements [47] on this material yield a centre to edge resistance of 325k at 700 C and 4k at 750 C. From top to bottom the corresponding resistances are 0.46 to 0.03 between 700 C and 750 C respectively. For typical values of photocurrent generated in the XPS spectrometer, significant charging shift would only be observed for a resistance of greater than 100M. The initial large shift in binding energy in the C 1s and O 1s, at temperatures up to 300 C, suggests that the presence of contaminants affects the electronic nature of the measured sample surface changing the band bending or band gap of the semiconducting SU-8, confirming that a true measurement of SU-8 is not achieved until 300 C.
@@
The C1 peak exhibits an increase in asymmetry during in-situ heating to higher pyrolysis temperatures. Figure 7 shows the variation of the asymmetry factor (, as defined by the Voigt fit) with pyrolysis temperature. A sharp increase in asymmetry occurs at 550 C that continues to a value of 0.95 at a pyrolysis temperature of 1000 C.
@@
Asymmetric peaks are due to a variety of factors such as vibrational modes, multi electron excitations and electron - hole pair generation. In carbon structures such as graphite, the asymmetry can be ascribed to electrons excited into the conduction band during the photoemission process, causing a loss of energy for the core level photoelectrons. As the sample becomes more conductive, the change in asymmetry can be correlated with changing conductivity [48, 49, 50].
@@
Figure 10A shows in more detail the valence band maximum of SU-8 as a function of temperature with spectra at selected temperatures shown in Figure 10B with a focus on photoelectrons emitted from near the Fermi Energy at 21.2 eV. The spectra were recorded with a small negative sample bias to enable measurement of the SECO, this value has been subtracted from the UPS spectra shown in Figures 9 and 10. As the temperature increases, a broad peak appears at 17 eV, that is most prominent at 400 C and disappears in the temperature range of 700 - 750 C. This coincides with the disappearance of O2 beyond 300 C (Figure 3) and the saturation of C-C peak position and is thus likely to be due to surface oxygen species (Figure 4). As the temperature approaches 1000 C, the VBM shows increasing emission at the Fermi energy (zero binding energy) at 21.2 eV (inset Figure 10B) that indicates the full transformation of insulating SU-8 to fully conducting glass-like carbon reported in the literature [15].
@@ Acknowledgements
Data availability
Information on the data underpinning this publication, including access details, can be found in the Cardiff University Research Data Repository at https://doi.org/10.17035/cardiff.29931761.v1